# Variational Rank Reduction Autoencoders

### v1

@@ 3.2 Real World data
We compare VRRAEs to the other proposed models on three real-world datasets: the MNIST, CIFAR-10, and CelebA. To evaluate their performance, autoencoders’ test reconstruction error, as well as the Fréchet Inception Distance (FID [^2] More details about how the FID was computed can be found in Appendix A.4.) score for both interpolation and random generation, are reported. For interpolation, we randomly sample $250$ couples of images and interpolate linearly in the latent space to get 5 new generated images on which the FID score is computed. Random generation is performed by using a Gaussian Mixture Model on the latent space and randomly sampling 1250 images from it. The results, documented in Table 2, show that VRRAEs are able to achieve a better performance than both RRAEs and VAEs on almost all the datasets presented. First, we note that the reconstruction error on the test set is always smaller for VRRAEs. This is mainly due to the fact that the KL-divergence isn’t as prominent as in VAEs since there is an existing regularization in RRAEs. The lower reconstruction error, together with the regularization from both the KL divergence and the truncated SVD, allows VRRAEs to achieve the best FID scores on almost all the datasets. More details about the hyperparameters and how fair comparative training was achieved for all autoencoders can be found in Appendix A.3.
Table 2: Quantitative comparison across datasets. The third column for each dataset represents the test reconstruction error, while the first two columns document the FID score for interpolation and random generation respectively.
@@
We present some examples of interpolated images on the CelebA dataset in Figure 4. We also show some generated MNIST and CelebA samples in Figures 5, and 6. More interpolated/generated images of all three datasets can be found in Appendix A.5. Overall the results show that VRRAEs outperform both deterministic RRAEs and VAEs, for the exception of the MNIST random generation where the SVD regularization doesn’t seem to be as important.
@@ 4 Ablation Study
    RRAE + VAE: A naive combination between an RRAE and a VAE, with $f$ being a trainable linear map instead of the identity. More details behind why this architecture is expected to perform worse can be found in Appendix A.1.
@@
A can be seen in both reconstruction errors and FIDs for interpolation/random generation, setting $f=I$ does not by itself bring any advantages when implemented on VAEs. However, it is necessary to get a good performance with VRRAEs. Further, the additional regularization caused by the KL divergence is necessary in most of the problems (except on CIFAR-10) to get a meaningful latent space (more details about the effect of the KL divergence can be found in Appendix A.2).
@@ 5 Conclusion
All in all, we presented in this paper a novel architecture, Variational Rank Reduction Autoencoders. The main objective behind VRRAEs is to learn a probabilistic model to be able to generate new samples (similarly to a VAE), while benefiting from the regularization imposed by RRAEs. Our results show that by preserving the deterministic values as the mean values of the distribution (i.e. $f=I$), VRRAEs are better regularized than VAEs, and are more robust to posterior collapse. They also outperform significantly their deterministic version RRAEs. Empirically the robustness to collapse was showcased on a synthetic dataset of small size, while the regularization effect was shown on three real-world datasets; MNIST, CelebA, and CIFAR-10, over which VRRAEs achieved better FID scores, over almost all generation/interpolation tasks compared to VAEs and RRAEs.
Acknowledgments and Disclosure of Funding
We thank SKF Magnetic Mechatronics for funding the research. We also thank the Google TPU Research cloud (TRC) program for giving us the resources needed to run all of the experiments.
@@ A.1 Naive RRAE + VAE
In this part, we present a naive combination of the concepts of an RRAE and a VAE, with $f=NN$, being any learnable function. While this approach makes RRAEs variational, the sampled coefficients $\tilde{\alpha}$ are problematic for two main reasons:
@@
    which means $\tilde{s}_{i}=\|\tilde{\alpha}_{i}^{T}\|_{2}$, and $\tilde{V}_{i}^{T}=\frac{\tilde{\alpha}_{i}^{T}}{\|\tilde{\alpha}_{i}^{T}\|_{2}}$ are the sampled singular values/vectors respectively. However, by choosing $f$ and $g$ to be Linear maps, we can not enforce that $\|\tilde{\alpha}_{i}^{T}\|_{2}\geq\|\tilde{\alpha}_{j}^{T}\|_{2},\forall i<j$. Accordingly, nothing guarantees that $\tilde{s}_{i}\geq\tilde{s}_{j},\forall i<j$, which means that the sampled singular values are not sorted. This complicates the training and reduces its stability, and complicates the proposal of an adaptive algorithm like the one proposed in [21].
@@
Note that empirically, we showed that choosing $f$ to be any learnable linear map does not achieve better results in the ablation study in Section 4.
@@ A.2 The KL divergence in VRRAEs
$\text{sum}(\bar{\alpha}_{\mu}\odot\bar{\alpha}_{\mu})=\text{sum}\left(\begin{bmatrix}&\text{sum}\left(\bar{s}_{1}^{2}\left(\bar{V}_{1}^{T}\odot\bar{V}_{1}^{T}\right)\right)&\\ &\vdots&\\ &\text{sum}\left(\bar{s}_{k^{*}}^{2}\left(\bar{V}_{k^{*}}^{T}\odot\bar{V}_{k^{*}}^{T}\right)\right)&\end{bmatrix}\right)=\text{sum}\left(\begin{bmatrix}&\bar{s}_{1}^{2}\,\,\,\text{sum}\left(\bar{V}_{1}^{T}\odot\bar{V}_{1}^{T}\right)&\\ &\vdots&\\ &\bar{s}_{k^{*}}^{2}\,\,\,\text{sum}\left(\bar{V}_{k^{*}}^{T}\odot\bar{V}_{k^{*}}^{T}\right)&\end{bmatrix}\right),$
@@ A.3 Model Architecture and Training Parameters
A.4 FID computation
@@
    The FID is computed as presented for instance in [17] between the feature vectors of the training data, and those of the generated images.
A.5 More Interpolations
Since CIFAR-10 images are of low resolution ($28\times 28$), and the autoencoders compress these images, the resulting photos are of low quality. An example of interpolated images of the CIFAR-10 dataset can be seen in Figure A.5-3,
@@ A.5 More Interpolations
We also present three more interpolation examples on the CelebA dataset in Figures A.5-4, A.5-5, and A.5-6. An interpolation example for the MNIST can be found in Figure A.5-7.
@@
As can be seen in all of the figures above, VRRAEs have the sharpest reconstruction and the best interpolation compared to both VAEs and RRAEs.

### v2

@@ 4.2 Real World data
We compare VRRAEs to the other proposed models on three real-world datasets: the MNIST, CIFAR-10, and CelebA. To evaluate their performance, autoencoders’ test reconstruction error, as well as the Fréchet Inception Distance (FID [^2] More details about how the FID was computed can be found in A.7.) score for both interpolation and random generation, are reported. For interpolation, we randomly sample $250$ couples of images and interpolate linearly in the latent space to get 5 new generated images on which the FID score is computed. Random generation is performed by using a Gaussian Mixture Model on the latent space and randomly sampling 1250 images from it. The results, documented in Table 2, show that VRRAEs are able to achieve a better performance than both RRAEs and VAEs on almost all the datasets presented. First, we note that the reconstruction error on the test set is always smaller for VRRAEs. This is mainly due to the fact that the KL-divergence isn’t as prominent as in VAEs since there is an existing regularization in RRAEs. The lower reconstruction error, together with the regularization from both the KL divergence and the truncated SVD, allows VRRAEs to achieve the best FID scores on almost all the datasets. More details about the hyperparameters and how fair comparative training was achieved for all autoencoders can be found in A.3.
Table 2: Quantitative comparison across datasets. The third column for each dataset represents the test reconstruction error, while the first two columns document the mean FID score (over 5 random seeds) for interpolation and random generation respectively. Standard deviations have been omitted since they’re of small magnitudes.
@@
We present some examples of interpolated images on the CelebA dataset in Figure 4. We also show some generated MNIST and CelebA samples in Figures 5, and 6. More interpolated/generated images of all three datasets can be found in A.8. Overall the results show that VRRAEs outperform both deterministic RRAEs and VAEs, for the exception of the MNIST random generation where the SVD regularization doesn’t seem to be as important.
@@ 5 Ablation Study
    RRAE + VAE: A naive combination between an RRAE and a VAE, with $f$ being a trainable linear map instead of the identity. More details behind why this architecture is expected to perform worse can be found in A.1.
@@
A can be seen in both reconstruction errors and FIDs for interpolation/random generation, setting $f=I$ does not by itself bring any advantages when implemented on VAEs. However, it is necessary to get a good performance with VRRAEs. Further, the additional regularization caused by the KL divergence is necessary in most of the problems (except on CIFAR-10) to get a meaningful latent space (more details about the effect of the KL divergence can be found in A.2).
@@ 6 Conclusion
All in all, we presented in this paper a novel architecture, Variational Rank Reduction Autoencoders. The main objective behind VRRAEs is to learn a probabilistic model to be able to generate new samples (similarly to a VAE), while benefiting from the regularization imposed by RRAEs. Our results show that by preserving the deterministic values as the mean values of the distribution (i.e., $f=I$), VRRAEs are better regularized than VAEs, and are more robust to posterior collapse. They also outperform significantly their deterministic version RRAEs. Empirically, the robustness to collapse was showcased on a synthetic dataset of small size, while the regularization effect was shown on three real-world datasets; MNIST, CelebA, and CIFAR-10, over which VRRAEs achieved better FID scores, over almost all generation/interpolation tasks compared to VAEs and RRAEs.
We thank SKF Magnetic Mechatronics for funding the research. We also thank the Google TPU Research cloud (TRC) program for giving us the resources needed to run all of the experiments. This work was also supported by Ministerio de Asuntos Económicos y Transformación Digital, Gobierno de España (Grant No. TSI-100930-2023-1) and Ministerio de Ciencia, Innovación y Universidades (Grant No. PID2023-147373OB-I00).
@@ A.1 Naive RRAE + VAE
In this part, we present a naive combination of the concepts of an RRAE and a VAE, with both $f$ and $g$ being learnable linear maps. While this approach makes RRAEs variational, the sampled coefficients $\tilde{\alpha}$ are problematic for two main reasons:
@@
    which means $\tilde{s}_{i}=\|\tilde{\alpha}_{i}^{T}\|_{2}$, and $\tilde{V}_{i}^{T}=\frac{\tilde{\alpha}_{i}^{T}}{\|\tilde{\alpha}_{i}^{T}\|_{2}}$ are the sampled singular values/vectors respectively. However, by choosing $f$ and $g$ to be linear maps, we can not enforce that $\|\tilde{\alpha}_{i}^{T}\|_{2}\geq\|\tilde{\alpha}_{j}^{T}\|_{2},\forall i<j$. Accordingly, nothing guarantees that $\tilde{s}_{i}\geq\tilde{s}_{j},\forall i<j$, which means that the sampled singular values are not sorted. This complicates the training and reduces its stability. Further, it makes the proposal of an adaptive algorithm like the one proposed in [33] impossible.
@@
Note that empirically, we showed that choosing $f$ to be a learnable linear map achieves worse results in the ablation study in Section 5.
@@ A.2 The KL divergence in VRRAEs
[…]
@@ A.3 Model Architecture and Training Parameters
A.4 Time complexity
In this section, we study the added time complexity in VRRAEs. Note that the main difference between both VAEs and VRRAEs is the SVD computation in the latent space. However, the latent matrix $Y$ is of size $(L\times\text{bs})$, where $L$ is the chosen latent space dimension, and bs is the batch size. Accordingly, the number of floating point operations (flops) for an SVD in a forward pass is:
$\text{n}_{\text{flops}}=\mathcal{O}\,(L\times\text{bs}\times\min(\text{bs},L))$
On the other hand, note that the backward pass through the SVD is only a matrix multiplication (as shown in the Appendix of [33]). In addition, the matrices being multiplied during the backward passare all of dimensions smaller than $\max(L,\text{bs})$. Consequently, the additional number of flops during backward propagation is around the following,
$\text{n}_{\text{back}}=\mathcal{O}\,\max(L,\text{bs})^{2}$
From the above, we can conclude that the additional time complexity of VRRAEs is similar to adding an additional layer of size $\max(L,\text{bs})$ to the network. While the overhead is not negligible, it can be considered small compared to the time complexity of the convolutional layers in the encoder/decoder. Empirically, both the required time for one forward/backward pass, as well as the total training time, can be found in Table A.4-1,
Table A.4-1: Training times for one forward/backward pass ($t$), and the total training time ($T$) for both VAEs and VRRAEs on all real-world datasets.
Model $t_{\text{MNIST}}$ $T_{\text{MNIST}}$ $t_{\text{CIFAR}}$ $T_{\text{CIFAR}}$ $t_{\text{CelebA}}$ $T_{\text{CelebA}}$
VAE 2.64 s 92.4 mn 0.315 s 20.47 mn 3.18 s 116.6 mn
VRRAE 2.79 s 97.7 mn 0.319 s 20.74 mn 3.39 s 124.3 mn
A.5 Effect of the batch size on the SVD
Since the SVD is computed on the latent space $Y$, which is of size $(L\times\text{bs})$, where $L$ is the latent space dimension, and bs is the batch size, the batch size has an effect on the SVD computation. In what follows, we show empirically that the batch size does not have a significant effect on the performance of VRRAEs. To do so, we trained VRRAEs on the synthetic dataset (i.e. the 2D Gaussian curves) for different batch sizes, and we documented the test reconstruction error in Table A.4-2.
Table A.4-2: Test Reconstruction Errors (in %) for different batch sizes on the synthetic dataset (i.e., 2D gaussians).
Batch size 8 16 32 64
Test error $10.03\pm 8.96$ $9.45\pm 2.13$ $8.47\pm 3.6$ $12.07\pm 4.1$
Note that the mean values are similar. On the other hand, the standard deviation is higher for a very small batch size. However, this is expected behavior when training Neural Networks, so further study is needed to conclude wether the difference in standard deviation is due to the SVD or not when the batch size is as small as $8$.
A.6 Comparison to other Autoencoders
In this section, we compare VRRAEs to other Autoencoder architectures on both MNIST and CelebA datasets. The FID values for both interpolation and random generation are taken from [33]. The results can be found in Table A.4-3.
Table A.4-3: Quantitative comparison across datasets. The third column for each dataset represents the test reconstruction error, while the first two columns document the mean FID score (over 5 random seeds) for interpolation and random generation respectively. Standard deviations have been omitted since they’re of small magnitudes.
MNIST CelebA
Model Inter. Rand. Rec. Inter. Rand. Rec.
Long 10.16 86.5 35.21 67.13 16.53 17.23
IRMAE 8.09 42.58 23.2 43.62 15.79 15.47
Cont 30.5 90.5 27.51 55.32 16.74 17.36
Sparse 6.1 57.67 61.32 87.2 16.91 16.57
AE 7.31 40.12 27.31 15.24 15.88 18.82
VAE 11.95 30.63 32.602 8.94 9.66 16.55
RRAE 6.68 45.46 27.90 13.27 13.70 17.52
VRRAE 5.89 38.77 26.00 7.06 7.60 15.03
A.7 FID computation
@@
    The FID is computed between the feature vectors of the training data, and those of the generated images.
A.8 More Interpolations
Since CIFAR-10 images are of low resolution ($28\times 28$), and the autoencoders compress these images, the resulting photos are of low quality. An example of interpolated images of the CIFAR-10 dataset can be seen in Figure A.5-3, more images can be found in Figures A.5-4, A.5-5, and A.5-7, where VRRAEs have the sharpest reconstruction and the best interpolation compared to both VAEs and RRAEs. Since CIFAR-10 images are of low resolution ($28\times 28$), and the autoencoders compress these images, the resulting photos are of low quality. An example of interpolated images of the CIFAR-10 dataset can be seen in Figure A.5-3, more images can be found in Figures A.5-4, A.5-5, and A.5-7, where VRRAEs have the sharpest reconstruction and the best interpolation compared to both VAEs and RRAEs.